Data were gathered from a simple random sample of cities. The variables are Violent Crime (crimes per 100,000 population), Police Officer Wage (mean S/hr), and Graduation Rate (%). The accompanying table shows a multiple regression model for the variables. Complete parts (a) through (d) EEB Click to view the regression table a) Write the regression model Police Officer Wage + Graduation Rate Violent Crime (Round to two decimal places as needed.) b) What does the coefficient of Police Officer Wage mean in the context of this regression model? an increase in equal to the coeficient V each dollar per hour of average wage increase Among cities with similar of Police Officer Wage c) In a city in which the average police officer wage is $21/hour and the high school graduation rate is 70%, what does this model estimate the violent crime rate would be? rimes per 100,000 population (Type an integer or a decimal rounded to two decimal places as needed.) d) Is this likely to be a good prediction? OA. No. The R2 value is fairly low so the prediction could be inaccurate O B. Yes. The regression assumptions are satisfied so the prediction will be very accurateData were gathered from a simple random sample of cities. The variables are Violent Crime (crimes per 100,000 population), Police Officer Wage (mean $/hr), and Graduation Rate (%). The accompanying table shows a multiple regression model for the variables. Complete parts (a) through (d).
Answer #1

(a) Regression Model

y^ = 1387.53 + 9.19 * Police officer wage - 16.23 * Graduation Rate

(b) Among cities with similar "Graduation Rate" each dollar per hour of average wage increases in "Police Officer Wage" , an increase in "Violent Crime" equal to coefficient of Police Officer wage.

(c) Here police officer wage = $ 21 /Hour

Hifgh School Graduation Rate = 70%

Violent Crime ^ = 1387.53 + 9.19 * 21 -16.23 * 70 = 444.42 crimes per 100,000 population

(d) Here as p - value is less than 0.001 so we will see that Yes, the regression are satisfied so the prediction will be accurate.